Alternative X has a first cost of $20000, an operating cost of $9000 per year, and a $5000 salvage value after 8 years. Alternative Y will cost $12,183 with an operating cost of $7,627 per year and a salvage value of $7,104 after 4 years. At an MARR of 0.12% per year, find the PW of machine Y?
Answer #1

PW (Present worth of Y) = -Initial cost - operating cost (P/A, i%, n) + salvage value (P/F. i%, n)

= -12183 - 7627(P/A, 0.12%, 4) + 7104(P/F, 0.12%, 4)

= -12183 - 7627*3.98803 + 7104*0.99521

= -33,530

Note:

I believe MARR is not 0.12% but 12%. In this case PW of Y will be -12183 - 7627(P/A, 12%, 4) + 7104(P/F, 12%, 4) = -30,834